# Retrieval-Augmented Generation for AI-Generated Content: A Survey

### v5

@@ IV-C3 Table for Question Answering
Tables can be integrated into prompts for query-based RAG. Both T-RAG [276] and OmniTab [277] concatenates the retrieved tables with the query to generate the answer. CARP [278] extracts hybrid chain of retrieved tables and passages for prompt construction. StructGPT [253] retrieves from multiple sources including KGs, tables, and databases. cTBLS [279] forms prompts with ranked tables after retrieval. Min et al. [280] integrated tabular data through table-to-text techniques, then experiments on both finetuning and RAG.
@@ IV-C4 Others
Prototype-KRG [281] integrates retrieved knowledge facts and dialogue prototypes into a GRU model through both hidden states and logits. SURGE [282] combines relevant subgraphs into the input for dialogue generation. RHO [283] fuses KG embedding of relevant entities and relations into textual embeddings during dialogue generation. K-LaMP [284] retrieves entities in history queries to construct prompt for query suggestion. ReSKGC [111] retrieves relevant triplets to complete triplet using Fid. G-Retriever [285] retrieves nodes and edges from textual graphs to construct subgraph and perform graph prompt tuning for QA.
@@ IV-D1 Image Generation
The retrieval process can not only help yield high-quality images even for rare or unseen subjects, but also reduces the parameter count and computational expense [45, 94, 103, 104, 102, 286, 105, 106]. For GAN-based model, RetrieveGAN [45] uses a differentiable retriever for image patch selection, facilitating end-to-end training. IC-GAN [94] models data as conditional distributions around each training instance, conditioning both the generator and discriminator on these instances. Recently, diffusion models beat GANs on image generation [287]. KNN-Diffusion [103] and RDM [104] train diffusion models conditioned on CLIP embeddings and image neighbors, enabling post-hoc conditioning on labels, prompts, and zero-shot stylization [105]. Beyond only images, Re-imagen [102] extends retrieval to image-text pairs for text-to-image generation, with interleaved guidance to balance the alignment between prompts and retrieval conditions. Retrieve&Fuse [286] prevents information loss of CLIP embeddings by concatenating retrieved and noised images before each U-Net attention block, allowing fully interaction via self-attention. RPG [288] retrieves representative images to construct in-context examples, and utilizes chain-of-thought reasoning [289] to plan out complementary subregions for compositional text-to-image diffusion.
@@ IV-D2 Image Captioning
Retrieval-augmented image captioning typically synthesises description with a collection of retrieved captions. MA [120] augments via a memory bank, built with historical context and target word of image-text training set, and queried with inference context. Vocabulary distribution is computed based on retrieved entries, and interpolated with the original prediction. In adversarial training, RAMP [290] takes retrieved captions as discriminator reference, and employs memory-augmented attention and copying mechanisms for better utilization of retrieved captions.
@@ IV-H2 Biomedical Informatics Enhancement
Several recent studies have improved the expressiveness of LLM by retrieving information from biomedical domain-specific databases, thereby augmenting the model’s capabilities to provide valuable guidance for tasks in the medical field. PoET [312] is an autoregressive generative model based on a variant of Transformer that integrates a retrieval mechanism to enable prompt augmentation, thereby expediting the prediction of fitness properties for protein variants. Chat-Orthopedist [93] enhances ChatGPT with a retrieval-augmented mechanism focused on adolescent idiopathic scoliosis (AIS), utilizing an external knowledge base for precise responses. BIOREADER [313] is the first retrieval-enhanced text-to-text transformer-based model for biomedical natural language processing, incorporating the retrieved literature evidence into the model using a chunked-cross attention mechanism. MedWriter [314] employs a hierarchical retrieval-augmented generation method that combines report-level and sentence-level templates to produce coherent and clinically accurate medical reports from images. QA-RAG [315] employs a dual-track RAG strategy to enhance pharmaceutical compliance by effectively retrieving and integrating regulatory guidelines based on language model responses and user queries.
@@ IV-H3 Math Applications
Retrieval-augmented generation technology in mathematics streamlines problem-solving, boosts research innovation, and refines educational strategies. LeanDojo [316] boosts theorem proving by using retrieval-augmented methods to choose relevant premises from extensive mathematical libraries, improving automation and theorem generalization. RAG-for-math-QA [317] improves math question-answering by integrating a high-quality math textbook with retrieval-augmented generation, enhancing LLM-generated responses for middle-school algebra and geometry.
@@ V Benchmark
Another three benchmark, RAGAS [319], ARES [320] and TruLens [321], consider another three aspects with the help of a separate evaluator LLM: (1) Faithfulness, which examines factual errors in results when correct answers can be inferred from the retrieved contents; (2) Answer Relevance, which measures whether generated results actually address the queries; (3) Context Relevance, which judges whether retrieved content provides relevant knowledge while minimizing irrelevant information.
CRUD-RAG [322] categorizes RAG tasks into four types: Create, Read, Update, and Delete. It evaluates each category using text continuation, question answering (including single- and multi-document questions), hallucination modification, and open-domain multi-document summarization. MIRAGE [323] assesses RAG in the medical domain, focusing on the performance of medical question-answering systems. KILT [324] aligns Wikipedia snapshots to verify information accuracy, using BLEU scores to pinpoint relevant texts and filtering to uphold quality, thus providing diverse retrieval systems for evidence-backed predictions or citations.
@@ VI-B2 Flexible RAG Pipelines
RAG systems are progressively embracing flexible pipelines, such as recursive, adaptive, and iterative RAG. Alongside proper tuning and careful engineering of each component, the unique combination of retrieval sources, retrievers, generators, and RAG subsystems holds promise for handling complex tasks and enhancing overall performance. We eagerly anticipate pioneering exploration that will drive the evolution of even more innovative RAG systems.
@@ VI-B6 Combined with Other Techniques
RAG is orthogonal to other techniques that also aim to improve AIGC effectiveness, such as fine-tuning, reinforcement learning, chain-of-thought, and agent-based generation. The combining of these methods is still in its early stages, calling for further research to fully exploit their potential through novel algorithm designs. It is worthy to note that a recent notion appears “long-context models like Gemini 1.5 will replace RAG”. Nevertheless, this assertion overlooks RAG’s flexibility in managing dynamic information, encompassing both up-to-date and long-tail knowledge [334]. We expect RAG to benefit from long context generation, rather than being replaced by it.
@@ VII Conclusion
In this paper, we conducted a thorough and comprehensive survey on RAG within the context of AIGC, with a particular focus on augmentation foundations, enhancements, and applications. We first systematically organized and summarize the foundation paradigms in RAG, providing insights into the interaction between retrievers and generators. Then, we reviewed the enhancements that further improve the effectiveness of RAG, including the enhancements on each component or the entire pipeline. To facilitate researchers across diverse domains, we showcased practical applications of RAG in a range of modalities and tasks. Finally, we also presented existing benchmarks for RAG, discussed current limitations of RAG, and shed light on promising future directions.

### v6

@@ IV-C2 Knowledge-augmented Open-domain Question Answering
GNN-RAG [283] fuses LLMs’ language understanding with GNN’s reasoning prowess and employs a retrieval augmentation strategy to enhance KGQA performance.
@@ IV-C3 Table for Question Answering
Tables can be integrated into prompts for query-based RAG. Both T-RAG [290] and OmniTab [291] concatenates the retrieved tables with the query to generate the answer. CARP [292] extracts hybrid chain of retrieved tables and passages for prompt construction. StructGPT [266] retrieves from multiple sources including KGs, tables, and databases. cTBLS [293] forms prompts with ranked tables after retrieval. Min et al. [294] integrated tabular data through table-to-text techniques, then experiments on both finetuning and RAG. ERATTA [295] generates SQL code to extract table information, integrating it into the prompt to minimize model hallucination.
@@ IV-C4 Others
Prototype-KRG [296] integrates retrieved knowledge facts and dialogue prototypes into a GRU model through both hidden states and logits. SURGE [297] combines relevant subgraphs into the input for dialogue generation. RHO [298] fuses KG embedding of relevant entities and relations into textual embeddings during dialogue generation. K-LaMP [299] retrieves entities in history queries to construct prompt for query suggestion. ReSKGC [112] retrieves relevant triplets to complete triplet using Fid. G-Retriever [300] retrieves nodes and edges from textual graphs to construct subgraph and perform graph prompt tuning for QA. Hussien et al. [301] fuse the reasoning power of KG with the expressiveness of LLMs through RAG techniques. HippoRAG [302] excels in multi-hop question answering by emulating mammalian brain knowledge storage with KG triples and employing a personalized PageRank algorithm for retrieval.
@@ IV-D1 Image Generation
The retrieval process can not only help yield high-quality images even for rare or unseen subjects, but also reduces the parameter count and computational expense [45, 95, 104, 105, 103, 303, 106, 107]. For GAN-based model, RetrieveGAN [45] uses a differentiable retriever for image patch selection, facilitating end-to-end training. IC-GAN [95] models data as conditional distributions around each training instance, conditioning both the generator and discriminator on these instances.
Recently, diffusion models beat GANs on image generation [304]. KNN-Diffusion [104] and RDM [105] train diffusion models conditioned on CLIP embeddings and image neighbors, enabling post-hoc conditioning on labels, prompts, and zero-shot stylization [106]. Beyond only images, Re-imagen [103] extends retrieval to image-text pairs for text-to-image generation, with interleaved guidance to balance the alignment between prompts and retrieval conditions. Retrieve&Fuse [303] prevents information loss of CLIP embeddings by concatenating retrieved and noised images before each U-Net attention block, allowing fully interaction via self-attention. RPG [305] retrieves representative images to construct in-context examples, and utilizes chain-of-thought reasoning [306] to plan out complementary subregions for compositional text-to-image diffusion.
@@ IV-D2 Image Captioning
Retrieval-augmented image captioning typically synthesises description with a collection of retrieved captions. MA [121] augments via a memory bank, built with historical context and target word of image-text training set, and queried with inference context. In adversarial training, RAMP [307] takes retrieved captions as discriminator reference, and employs memory-augmented attention and copying mechanisms for better utilization of retrieved captions.
@@ IV-H2 Biomedical Informatics Enhancement
Several recent studies have improved the expressiveness of LLM by retrieving information from biomedical domain-specific databases, thereby augmenting the model’s capabilities to provide valuable guidance for tasks in the medical field.
PoET [329] is an autoregressive model using a transformer variant with a retrieval mechanism for prompt augmentation, speeding up the prediction of protein variant fitness properties. Chat-Orthopedist [94] enhances ChatGPT with a retrieval-augmented mechanism focused on adolescent idiopathic scoliosis (AIS), utilizing an external knowledge base for precise responses. BIOREADER [330] is the first retrieval-enhanced text-to-text transformer-based model for biomedical natural language processing, incorporating the retrieved literature evidence into the model using a chunked-cross attention mechanism. MedWriter [331] employs a hierarchical retrieval-augmented generation method that combines report-level and sentence-level templates to produce coherent and clinically accurate medical reports from images. QA-RAG [332] employs a dual-track RAG strategy to enhance pharmaceutical compliance by effectively retrieving and integrating regulatory guidelines based on language model responses and user queries. RAG-RLRC-LaySum [333] leverages biomedical text knowledge for llms, employing reinforcement learning and re-ranking techniques to enhance content relevance and readability of the output.
@@ IV-H3 Math Applications
Retrieval-augmented generation technology in mathematics streamlines problem-solving, boosts research innovation, and refines educational strategies.
LeanDojo [334] boosts theorem proving by using retrieval-augmented methods to choose relevant premises from extensive mathematical libraries, improving automation and theorem generalization. RAG-for-math-QA [335] improves math question-answering by integrating a high-quality math textbook with RAG, enhancing LLM-generated responses for middle-school algebra and geometry.
@@ V Benchmark
Three other benchmarks, RAGAS [337], ARES [338], and TruLens [339], evaluate three aspects using a separate evaluator LLM: (1) Faithfulness, assessing factual accuracy based on retrieved content; (2) Answer Relevance, determining if results address the queries; (3) Context Relevance, evaluating the relevance of retrieved content and its conciseness.
CRUD-RAG [340] divides RAG tasks into four types: Create, Read, Update, and Delete, assessing them through text continuation, question answering, hallucination correction, and open-domain multi-document summary. MIRAGE [341] assesses RAG in the medical domain, focusing on the performance of medical question-answering systems. KILT [342] aligns Wikipedia snapshots to verify information accuracy, using BLEU scores to pinpoint relevant texts and filtering to uphold quality, thus providing diverse retrieval systems for evidence-backed predictions or citations.
@@ VI-B2 Flexible RAG Pipelines
RAG systems are progressively embracing flexible pipelines, such as recursive, adaptive, and iterative RAG. With precise tuning and meticulous engineering, the unique blend of retrieval sources, retrievers, generators, and RAG subsystems promises to tackle complex tasks and boost overall performance. We eagerly anticipate pioneering exploration that will drive the evolution of even more innovative RAG systems.
@@ VI-B6 Combined with Other Techniques
RAG is orthogonal to other techniques that also aim to improve AIGC effectiveness, such as fine-tuning, reinforcement learning, chain-of-thought, and agent-based generation. The combining of these methods [352] is still in its early stages, calling for further research to fully exploit their potential through novel algorithm designs. It is worthy to note that a recent notion appears “long-context models like Gemini 1.5 will replace RAG”. Nevertheless, this assertion overlooks RAG’s flexibility in managing dynamic information, encompassing both up-to-date and long-tail knowledge [353]. We expect RAG to benefit from long context generation, rather than being replaced by it.
@@ VII Conclusion
In this paper, we conducted a thorough and comprehensive survey on RAG within the context of AIGC, with a particular focus on augmentation foundations, enhancements, and applications. We first systematically organized and summarized the foundation paradigms in RAG, providing insights into the interaction between retrievers and generators. Then, we reviewed the enhancements that further improve the effectiveness of RAG, including the enhancements on each component or the entire pipeline. To facilitate researchers across diverse domains, we showcased practical applications of RAG in a range of modalities and tasks. Finally, we also presented existing benchmarks for RAG, discussed current limitations of RAG, and shed light on promising future directions.